For any nxn matrix A, use the SVD to show that there is an nxn orthogonal matrix Q such that AAT = QT(ATA)Q.
Answer #1

let A be the n \times n matrix of n-dimensional points

By the singular value decomposition:

A - \cup \times\cap\times V ^ T

here

\cup ( n\timesn ) is orthogonal \cup ^T \cup = I

\cap ( n\timesn) has r positive singular value is descending order units diagonal

V( n\timesn ) is orthogonal V^T V = I

columns of value the orthogonal eigenvector of A^T A

= ( \cup \times \cap \times V^T) ^T (\cup \times \cap\times V^T)

=  \cup \times \cap^T \times   \cap\times  V^T

=\cup \times \cap^2 \times V^T

columns of value the orthogonal eigenvector of A A^T

=(\cup \times \cap\times V^T) ( \cup \times \cap \times V^T) ^T

= \cup \times \cap \times  \cap^T\times  V^T

=\cup \times \cap^2 \times V^T

A^T A=(QR)^T (QR)

=R^TR [\because Q^TQ = I ]

we have , A A^T = Q(RR^T)Q^T  

so, if it  is an eigen vector for AA^T,then Q^TQ is an eigenvector for RRT.

Hence ,the matrix V becomes Q^T \cup .